a women has 3 times as many dimes as nickels. she has $4.20. How many of each cion does she have?
Answer by josmiceli(19441)

d = number of dimes
n = number of nickels
10d + 5n = 420
d = 3n
10*3n + 5n = 420
30n + 5n = 420
35n = 420
n = 12
d = 3(12)
d = 36
She has 12 nickels and 36 dimes
check
5(12) + 10(36) = 420
60 + 360 = 420
420 = 420
OK